OpenStudy (anonymous):

What are the steps to solving 2|x−5|+1≥3? The answers is x≥6 or x≤4.

OpenStudy (kc_kennylau):

1. take 1 from both sides 2. divide both sides by 2 3. square both sides 4. put all thing to the LHS 5. solve the quadratic equation

OpenStudy (anonymous):

Thanks appreciate the help.

OpenStudy (kc_kennylau):

thanks

OpenStudy (ranga):

2|x − 5| + 1 ≥ 3 2|x − 5| ≥ 2 |x - 5| >= 1 means x - 5 >= 1 or x - 5 <= -1 x >= 6 or x <= 4

OpenStudy (ranga):

If |y| >= a it means y >= a or y <= -a

OpenStudy (anonymous):

Makes a lot more sense thanks.

OpenStudy (ranga):

You are welcome. If |Z| <= C it means -C <= Z <= +C If |Z| >= C it means Z <= -C or Z >= +C